OpenStudy (countrygirl1431):

Please help Algebra 2

OpenStudy (countrygirl1431):

Perform the indicated operation.

OpenStudy (igreen):

Flip the 2nd fraction and multiply

OpenStudy (countrygirl1431):

@iGreen

OpenStudy (countrygirl1431):

\[\frac{ (a+3)^2 }{ a-3 }\times \frac{ a^2-9 }{5 }\]

OpenStudy (igreen):

Correct, and can you multiply those 2 together?

OpenStudy (igreen):

Multiply the numerators together and multiply the denominators together.

OpenStudy (countrygirl1431):

im not sure how

OpenStudy (countrygirl1431):

i know it starts off \[2a^2\]

OpenStudy (igreen):

Okay, for the numerator we have \(\sf (a + 3)(a + 3)(a^2 - 9)\)

OpenStudy (igreen):

Can you multiply \(\sf (a + 3)(a + 3)\)?

OpenStudy (countrygirl1431):

a+9

OpenStudy (igreen):

No.. \(\sf a \times a\) \(\sf a \times 3\) \(\sf 3 \times a\) \(\sf 3 \times 3\) Multiply those

OpenStudy (countrygirl1431):

im confused

OpenStudy (igreen):

\(\sf a \times a \rightarrow a^2\) \(\sf a \times 3 \rightarrow 3a\) \(\sf 3 \times a \rightarrow 3a\) \(\sf 3 \times 3 \rightarrow 9\)

OpenStudy (igreen):

Which gives us \(\sf a^2 + 6a + 9\), now multiply that to \(\sf a^2 - 9\)

OpenStudy (countrygirl1431):

ok i thought tht was it but i wasnt sure

OpenStudy (countrygirl1431):

so would this be a^4

OpenStudy (countrygirl1431):

or 6a^5

OpenStudy (igreen):

You gotta multiply all of them

OpenStudy (igreen):

\(\sf a^2 \times a^2\) \(\sf a^2 \times -9\) \(\sf 6a \times a^2\) \(\sf 6a \times -9\) \(\sf 9 \times a^2\) \(\sf 9 \times -9\)

OpenStudy (igreen):

The first one equals \(\sf a^4\), that's correct.

OpenStudy (countrygirl1431):

the 2nd 1 is a^2-9

OpenStudy (igreen):

No, it's \(\sf -9a^2\)

OpenStudy (countrygirl1431):

oh ok

OpenStudy (countrygirl1431):

3=6a^3

OpenStudy (igreen):

Correct.

OpenStudy (countrygirl1431):

would 4 be -54a

OpenStudy (countrygirl1431):

5 wouldbt 9a^2 and 6 would be -81

OpenStudy (igreen):

4 is correct.

OpenStudy (igreen):

So are 5 and 6.

OpenStudy (countrygirl1431):

ok

OpenStudy (igreen):

That gives us: \(\sf a^4 - 9a^2 + 6a^3 - 54a + 9a^2 - 81\)

OpenStudy (igreen):

Combine like terms and you got: \(\sf a^4 + 6a^3 - 54a - 81\)

OpenStudy (igreen):

Now simplify the denominator. \(\sf 5(a - 3)\)

OpenStudy (countrygirl1431):

5a-3

OpenStudy (igreen):

Multiply 5 to both terms.

OpenStudy (igreen):

\(\sf 5 \times a\) \(\sf 5 \times -3\)

OpenStudy (countrygirl1431):

5a and -15

OpenStudy (igreen):

Yes, that gives us: \(\sf\dfrac{ a^4 + 6a^3 - 54a - 81}{5a - 15}

OpenStudy (countrygirl1431):

i cant read it @green_1

OpenStudy (countrygirl1431):

@iGreen

OpenStudy (countrygirl1431):

sorry @green_1 i meant to tag someone else

OpenStudy (green_1):

you mean the equation above

OpenStudy (countrygirl1431):

the last thing posted i cant read it

OpenStudy (countrygirl1431):

@iGreen could u repost the last thing